OpenStudy (anonymous):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)2. You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (anonymous):

This problem is asking you to make a mistake when using power rules. It wants you to first calculate \[(5-3)^2 = (2)^2 = 4\] and then incorrectly use the power rule described for a product of powers, raised to a power. That is, this problem is asking you to purposely make the mistake of doing this: \[(5-3)^2 = (5^2 - 3^2) = (25 - 9) = 16\] Clearly, you can see that 16 does not equal 4, so that was the incorrect way to work this problem.

OpenStudy (anonymous):

thank u!!

OpenStudy (anonymous):

You're welcome!